# Reducing Object Hallucination in Large Audio- Language Models via Audio-Aware Decoding

### v1

@@ Abstract
Large Audio-Language Models (LALMs) can take audio and text as the inputs and answer questions about the audio. While prior LALMs have shown strong performance on standard benchmarks, there has been alarming evidence that LALMs can hallucinate what is presented in the audio. To mitigate the hallucination of LALMs, we introduce Audio-Aware Decoding (AAD), a lightweight inference‐time strategy that uses contrastive decoding to compare the token prediction logits with and without the audio context. By contrastive decoding, AAD promotes the tokens whose probability increases when the audio is present. We conduct our experiment on object hallucination datasets with three LALMs and show that AAD improves the F1 score by 0.046 to 0.428. We also show that AAD can improve the accuracy on general audio QA datasets like Clotho-AQA by 5.4% to 10.3%. We conduct thorough ablation studies to understand the effectiveness of each component in AAD.
@@ II-B Audio-Aware Decoding (AAD)
We adjust the probability for each candidate token at step $t$ using the following contrastive decoding formulat [33, 37, 38]:
@@
where $\alpha$ is a hyperparameter that determines how much to emphasize the influence of the audio context versus the model’s prior knowledge. A higher $\alpha$ places more weight on the audio context and downweights contributions from the model’s prior knowledge withou hearing the audio. At each time step, the token is sampled based on Equation 2 and auto-regressively generates the output sequence.
@@ V-B Do We Need Prefix Prompts in AAD?
Model Acc F1 $\Delta$Acc(%)/$\Delta$F1(%) Random Sampling Qwen2‐Audio‐7B 0.656 0.555 -4.9/-7.5 SALMONN‐7B 0.577 0.328 -15.6/-33.3 SALMONN‐13B 0.706 0.604 -1.4/-1.7 Adversarial Sampling Qwen2‐Audio‐7B 0.469 0.338 +1.1/-0.9 SALMONN‐7B 0.502 0.276 -0.3/-14 SALMONN‐13B 0.524 0.393 +0.5/+7.0 Popular Sampling Qwen2‐Audio‐7B 0.484 0.374 -1.8/-7.0 SALMONN‐7B 0.532 0.360 -4.3/-18.0 SALMONN‐13B 0.586 0.516 +1.8/+7.7
@@ VII Conclusion, Limitations, and Future Work
While AAD is a promising method, there is an important limitation of this method: For one question, we need to forward the through the LALM twice, one with the audio and one without it. While this creates compute overhead compared to the default setting, the two forward passes in AAD can be done in parallel. As a result, AAD should incur little to no latency compared with the default setting. Based on the results in Table I, trading off the inference compute overhead with the performance gain seems to be a reasonable choice. Additionally, due to the nature of the datasets we use, we only evaluate object hallucination on yes/no questions. It will be interesting to explore other types of object hallucination of LALMs by constructing more diverse benchmarks. We leave this as future work.

### v2

@@ Abstract
Large Audio-Language Models (LALMs) can take audio and text as the inputs and answer questions about the audio. While prior LALMs have shown strong performance on standard benchmarks, there has been alarming evidence that LALMs can hallucinate what is presented in the audio. To mitigate the hallucination of LALMs, we introduce Audio-Aware Decoding (AAD), a lightweight inference‐time strategy that uses contrastive decoding to compare the token prediction logits with and without the audio context. By contrastive decoding, AAD promotes the tokens whose probability increases when the audio is present. We conduct our experiment on object hallucination datasets with three LALMs and show that AAD improves the F1 score by 0.046 to 0.428. We also show that AAD can improve the accuracy on general audio QA datasets like Clotho-AQA by 5.4% to 10.3%. We conduct thorough ablation studies to understand the effectiveness of each component in AAD. [^1] https://github.com/GillbertHsu/Audio-Aware-Decoding
@@ II-B Audio-Aware Decoding (AAD)
We adjust the probability for each candidate token at step $t$ using the following contrastive decoding formula [33, 37, 38]:
@@
where $\alpha$ is a hyperparameter that determines how much to emphasize the influence of the audio context versus the model’s prior knowledge. A higher $\alpha$ places more weight on the audio context and downweights contributions from the model’s prior knowledge without hearing the audio. At each time step, the token is sampled based on Equation 2 and auto-regressively generates the output sequence.
@@ V-B Do We Need Prefix Prompts in AAD?
[…]
@@ VII Conclusion, Limitations, and Future Work
While AAD is a promising method, there is an important limitation of this method: For one question, we need to forward the through the LALM twice, one with the audio and one without it. While this creates compute overhead compared to the default setting, the two forward passes in AAD can be done in parallel. As a result, AAD should incur little to no latency compared with the default setting. We tested the inference speed on 3,000 data samples with and without AAD. When computation is not parallelized, inference time increases by nearly 60%. However, when run in parallel, the increase is only around 19%. Based on the results in Table I, trading off the inference compute overhead with the performance gain seems to be a reasonable choice. Additionally, due to the nature of the datasets we use, we only evaluate object hallucination on yes/no questions. It will be interesting to explore other types of object hallucination of LALMs by constructing more diverse benchmarks. We leave this as future work.